Good
Please assist me with this question

Use the numerical method to calculate the BEP of a wine company that sells a 1000 litres of wine per week at 12 per litre.they produce wine at R8 per litre and their total fixed cost to operate their plant and to produce their wine is R2000 per week

Answer by Theo(13342)

revenue is equal to 1000 * 12.
cost is equal to 2000 + 1000 * 8.

not exactly sure what you're looking for.

based on the numbers you provided, they are making a profit.

revenue - cost = profit.

1000 * 12 - 1000 * 8 - 2000 = 1000 * (12-8) - 2000 = 1000 * 4 - 2000 = 4000 - 2000 = 2000.

if you are looking for the number of liters they can produce to break even, given that the price is 12 and the cost is 8, then the formula would be:

12 * x = 2000 + 8 * x, where x is the number of liters to produce.

subtract 8 * x from both sides of this equation to get 12 * x - 8 * x = 2000.

solve for x to get 4 * x = 1000 which leads to x = 500.

at the price of 12 and the cost of 8, they will break even if they make and sell 500 liters of wine.

revenue would be 500 * 12 = 6000
cost would be 2000 + 500 * 8 = 2000 + 4000 = 6000

here's a reference i found on the web about break even point analysis.

https://accountingexplained.com/managerial/cvp-analysis/break-even-point-equation-method

the formula is px = px + fc + profit

the break even point is when the profit is 0.

the formula then becomes px = vx + fc

p is the selling price per unit.
v is the variable cost per unit
x is the number of units
fc is the fixed cost

it's clear that, when they sell 1000 liters at 12 a liter and the cost is 8 a liter and the fixed cost is 2000, that they make a profit.

if you want to know what price they can sell the wine at in order to break even, then that's a different type of problem.

given that they sell 1000 liters a week, you can find the break even price by using the following formula.

1000 * x = 2000 + 1000 * 8

divide both sides of this equation by 1000 to get:

x = 2 + 8 which makes x = 10.

they could sell the wine at 10 each and break even.

1000 * 10 = 10,000
2000 + 1000 * 8 = 10,000.

so, .........

at the price of 12 each and the cost of 8 each plus 2000, they can break even if they sell 500 liters a week.

if they sell 1000 liters a week, they can break even if the price of each is 10 and the cost of each is 8 plus a fixed cost of 2000.